# Faster Quantum Simulation Of Markovian Open Quantum Systems Via Randomisation

### v2

@@ Abstract
When simulating the dynamics of open quantum systems with quantum computers, it is essential to accurately approximate the system’s behaviour while preserving the physicality of its evolution. Traditionally, for Markovian open quantum systems, this has been achieved using first and second-order Trotter-Suzuki product formulas or probabilistic algorithms. In this work, we introduce novel non-probabilistic algorithms for simulating Markovian open quantum systems using randomisation. Our methods, including first and second-order randomised Trotter-Suzuki formulas and the QDRIFT channel, not only maintain the physicality of the system’s evolution but also enhance the scalability and precision of quantum simulations. We derive error bounds and step count limits for these techniques, bypassing the need for the mixing lemma typically employed in Hamiltonian simulation proofs. We also present two implementation approaches for these randomised algorithms: classical sampling and quantum forking, demonstrating their gate complexity advantages over deterministic Trotter-Suzuki product formulas. This work is the first to apply randomisation techniques to the simulation of open quantum systems, highlighting their potential to enable faster and more accurate simulations.
@@ 1 Introduction
First and second-order TS product formulas have been utilised to simulate OQS [22], as they guarantee that the approximations are CPTP maps. However, higher-order TS product formulas are infeasible due to their recursive construction which results in non-CPTP maps. Alternative algorithms for simulating OQS have been proposed [23, 24, 25, 26], most of which are probabilistic, introducing a non-zero probability of failure. Despite their precision and gate complexity improvements, the quest for non-probabilistic algorithms for OQS simulation remains crucial.
@@
Inspired by the use of randomisation for Hamiltonian simulation [11, 9], we present two novel non-probabilistic methods for simulating Markovian OQS that make use of randomisation. Both methods maintain the physicality of the system’s evolution by producing CPTP approximations of the ideal evolution. The first method employs randomised TS product formulas up to the second order for the simulation of Markovian OQS. The randomised product formulas improve the scaling of the gate complexity with respect to the number of terms in the generator of the quantum channel, thereby facilitating faster simulation of Markovian OQS when compared to deterministic TS product formulas. The second method is inspired by the QDRIFT approach [9] for Hamiltonian simulation. In this case, the gate complexity of the QDRIFT-inspired method is independent of the number of terms in the generator, making it particularly suitable for large systems or those with numerous interacting components, such as the 2D dissipative Jaynes-Cummings model with neighbour-neighbour interaction [27]. For both the randomised product formulas method and the QDRIFT method, we derive error and number of step count bounds. These bounds do not rely on the mixing lemma by Campbell and Hastings [28, 29], which applies only to Hamiltonian simulation.
@@
To implement both the randomised product formulas and the QDRIFT method, we propose the use of Classical Sampling (CS) and Quantum Forking (QF). CS involves constructing a gate set by sampling from a discrete distribution on a classical computer. In contrast, QF [30] involves sampling directly on a quantum computer to implement the randomisation. We demonstrate that the gate complexities for the circuits constructed using both CS and QF are efficient relative to the number of terms in the generator and that they outperform the deterministic TS product formulas outlined in Section 2.
@@ 2.1 Background
for $k=2,...,M$. Sometimes it will be convenient to absorb the decay rates $\gamma_{k}$ into each $\mathcal{L}_{k}$, by defining ^$\hat{\mathcal{L}}_{k}=\gamma_{k}\mathcal{L}_{k}$ the generator can be written as
@@ 2.2 Deterministic Digital Simulation of Markovian Open Quantum Systems
The main goal of digital quantum simulation of Markovian OQS is to find novel ways of construct an approximation $\tilde{T}_{t}$ of the total evolution $T_{t}=\exp(t\mathcal{L})$ such that for a given GKSL generator $\mathcal{L}$, a precision $\epsilon>0$, a simulation time $t\geq 0$ and a distance measure dist(,)($\cdot$,$\cdot$)( ⋅ , ⋅ ),
@@
where $\tilde{T}_{t}$ could be implemented on a quantum computer efficiently.
The most common way to obtain this approximation is by using Trotter-Suzuki (TS) product formulas [5, 6]. Using TS product formulas, the total evolution $T_{t}=\exp(t\sum_{k=1}^{M}\hat{\mathcal{L}}_{k})$ is approximated as some product of simpler channels up to a precision $\epsilon\geq 0$ when using the diamond norm as the distance measure for the quantum channels. The same approach was introduced in [22], however, the $1\rightarrow 1$ Schatten norm was used instead of the diamond norm. We start by dividing the time $t\geq 0$ into $N\in\mathbb{N}$ steps so that we have a small time step $\tau=t/N$. Next we construct approximations of the total channel $T_{t}$ using TS product formulas that approximate the channel for a small time step $T_{\tau}$ and then simulate the TS product formula $N$ times. For example, we can approximate the evolution $T_{\tau}$ up to first order by using the following deterministic TS product formula,
@@
Let $N\in\mathbb{N}$ and $\Lambda:=\max_{k}\left\|\hat{\mathcal{L}}_{k}\right\|_{\diamond}$. Then,
$\left\|T_{t}-S_{1}^{(det)}\left(\frac{t}{N}\right)^{N}\right\|_{\diamond}\leq\epsilon,$ (21)
where,
$\epsilon\geq\frac{t^{2}\Lambda^{2}M^{2}}{N},$ $N\geq\frac{t^{2}\Lambda^{2}M^{2}}{\epsilon}.$ (22)
The proof of Theorem 1. can be found in Appendix B
@@
Let $N\in\mathbb{N}$ and $\Lambda:=\max_{k}\left\|\hat{\mathcal{L}}_{k}\right\|_{\diamond}$. Then,
$\left\|T_{t}-S_{2}^{(det)}\left(\frac{t}{N}\right)^{N}\right\|_{\diamond}\leq\epsilon,$ (24)
where,
$\epsilon\geq\frac{M^{3}t^{3}\Lambda^{3}}{3N^{2}},$ $N\geq\frac{M^{3/2}t^{3/2}\Lambda^{3/2}}{\sqrt{3\epsilon}}.$ (25)
The proof of Theorem 2. can be found in Appendix A.
Now that we have shown that we can approximate $T_{t}$ by TS product formulas, all that is left is to construct a quantum circuit that implements this product formula on a quantum computer. Since our product formula is a quantum channel, one needs to use a unitary dilation, for example the Stinespring representation of the channel [33, 34], to construct a quantum circuit. However, in this work, when we draw quantum circuits we will only show the action of the channel on the state as this keeps the diagrams concise. To make clear how the diagrams should be interpreted, we note that wires in our circuits correspond to density matrices of a subsystem and gates correspond to quantum channels, thereafter the usual rules of quantum circuits may be inferred. As an illustrative example consider the deterministic product formula $S_{2}^{(det)}(t/N)^{N}$ that approximates the total evolution $T_{t}$ up to a precision $\epsilon$. This means that the output of the quantum circuit that implements $S_{2}^{(det)}$ is a density matrix $\tilde{\rho}(t)$ that is a distance $\epsilon/2$ from the density matrix $\rho(t)$. One can easily see this by using the definition of the trace distance between states i.e. $d_{tr}(\rho(t),\tilde{\rho}(t))$ and inequality (15),
$d_{tr}(\rho(t),\tilde{\rho}(t))$ $=\frac{1}{2}\norm{\rho(t)-\tilde{\rho}(t)}_{1},$
$=\frac{1}{2}\norm{T_{t}(\rho(0))-S_{2}^{(det)}(t/N)^{N}\rho(0)}_{1},$
$\leq\frac{1}{2}\left\|T_{t}-S_{2}^{(det)}(t/N)^{N}\right\|_{\diamond},$
@@
[…]
@@ 3 First Order Randomised Trotter-Suzuki Formula
Given the generator $\mathcal{L}$, as in equation (10), of a quantum channel $T_{t}$ and some time $t\geq 0$. Define the first order randomised product formula as in equation (30). Let $N\in\mathbb{N}$ and $\Lambda:=\max_{k}\left\|\hat{\mathcal{L}}_{k}\right\|_{\diamond}$. Then,
$\left\|T_{t}-S_{1}^{(ran)}\left(\frac{t}{N}\right)^{N}\right\|_{\diamond}\leq\epsilon,$ (31)
where,
$\epsilon\geq\frac{M^{3}t^{3}\Lambda^{3}}{3N^{2}},$ (32)
$N\geq\frac{(t\Lambda M)^{3/2}}{\sqrt{3\epsilon}}.$ (33)
@@
For the second term in equation (3), we use the fact that the diamond norm is sub-multiplicative and ^$\left\|\hat{\mathcal{L}}_{i}\right\|_{\diamond}\leq\Lambda$ for all $i=1,...,M$ to show that
^$\left\|\hat{\mathcal{L}}_{1}^{j_{1}}\hat{\mathcal{L}}_{2}^{j_{2}}...\hat{\mathcal{L}}_{M}^{j_{M}}\right\|_{\diamond}\leq\left\|\hat{\mathcal{L}}_{1}\right\|_{\diamond}^{j_{1}}...\left\|\hat{\mathcal{L}}_{M}\right\|_{\diamond}^{j_{M}}\leq\Lambda^{j_{1}+...+j_{M}}.$ (52)
@@
Now using Lemma 1. from Appendix B, we can compute the restricted sums in equations (53) and (54). We then get the final bounds on these terms as

### v3

@@ Abstract
When simulating the dynamics of open quantum systems with quantum computers, it is essential to accurately approximate the system’s behaviour while preserving the physicality of its evolution. Traditionally, for Markovian open quantum systems, this has been achieved using first and second-order Trotter-Suzuki product formulas or probabilistic algorithms. In this work, we introduce novel non-probabilistic algorithms for simulating Markovian open quantum systems using randomisation. Our methods, including first and second-order randomised Trotter-Suzuki formulas and the QDRIFT channel, not only maintain the physicality of the system’s evolution but also enhance the scalability and precision of quantum simulations. We derive error bounds and step count limits for these techniques, bypassing the need for the mixing lemma typically employed in Hamiltonian simulation proofs. We also present two implementation approaches for these randomised algorithms: classical sampling and quantum forking, demonstrating their gate complexity advantages over deterministic Trotter-Suzuki product formulas. This work systematically extends powerful randomisation techniques from Hamiltonian simulation to the general setting of Markovian open quantum systems, highlighting their potential to enable faster and more accurate simulations.
@@ 1 Introduction
First and second-order TS product formulas have been utilised to simulate OQS [22], as they guarantee that the approximations are CPTP maps. However, higher-order TS product formulas are infeasible due to their recursive construction which results in non-CPTP maps. Alternative algorithms for simulating OQS have been proposed [23, 24, 25, 26], most of which are probabilistic, introducing a non-zero probability of failure. Despite their precision and gate complexity improvements, the quest for non-probabilistic algorithms for OQS simulation remains crucial. While advanced deterministic methods that achieve nearly linear scaling in simulation time have been developed [16, 18, 23, 24, 25, 26, 27, 28], they often rely on complex subroutines with significant overhead. This motivates the development of simpler, lower-overhead methods, such as the product-formula-based approaches we explore here.
@@
Inspired by the use of randomisation for Hamiltonian simulation [11, 9], we present two novel non-probabilistic methods for simulating Markovian OQS that make use of randomisation. Both methods maintain the physicality of the system’s evolution by producing CPTP approximations of the ideal evolution. The first method employs randomised TS product formulas up to the second order for the simulation of Markovian OQS. The randomised product formulas improve the scaling of the gate complexity with respect to the number of terms in the generator of the quantum channel, thereby facilitating faster simulation of Markovian OQS when compared to deterministic TS product formulas. The second method is inspired by the QDRIFT approach [9] for Hamiltonian simulation. In this case, the gate complexity of the QDRIFT-inspired method is independent of the number of terms in the generator, making it particularly suitable for large systems or those with numerous interacting components, such as the 2D dissipative Jaynes-Cummings model with neighbour-neighbour interaction [29]. For both the randomised product formulas method and the QDRIFT method, we derive error and number of step count bounds. These bounds do not rely on the mixing lemma by Campbell and Hastings [30, 31], which applies only to Hamiltonian simulation. Our results show in all cases that the randomised algorithms have error bounds with improved dependence on the number of terms $M$ in the GKSL generator.
@@
To implement both the randomised product formulas and the QDRIFT method, we propose the use of Classical Sampling (CS) and Quantum Forking (QF). CS involves constructing a gate set by sampling from a discrete distribution on a classical computer. In contrast, QF [32] involves sampling directly on a quantum computer to implement the randomisation.
To compare these methods, we define the gate complexity as the total number of ”simple channels” required to implement the simulation algorithm. A simple channel is considered a fundamental building block of the evolution, such as the exponential of a summand of the generator or a controlled-SWAP channel used in the quantum forking circuits. The table below summarizes the gate complexities derived in this work, demonstrating a clear advantage over deterministic approaches in their dependence on $M$.
Method Of Simulation Gate Complexity
Second Order TS Deterministic $O\left((t\Lambda)^{3/2}M^{5/2}/\sqrt{\epsilon}\right)$
First Order TS Randomised (CS) $O\left((t\Lambda)^{3/2}M^{5/2}/\sqrt{\epsilon}\right)$
QDRIFT (CS) $O\left((t\Gamma\Omega)^{2}/\epsilon\right)$
First Order TS Randomised (QF) $O\left((t\Lambda)^{3/2}M^{5/2}/\sqrt{\epsilon}\right)$
QDRIFT (QF) $O\left((t\Gamma\Omega)^{2}M/\epsilon\right)$
Composite Formulas [33] (Only apples to restricted class of Lindbladians) $O\left(t^{3/2}/\epsilon^{1/2}\right)$
Effective Hamiltonian Methods [28] $O(t\mathrm{poly}(\log(t/\epsilon)))$
Repeated Interactions [27] $O(t^{1+(1/2k)}\mathrm{poly}(\log(t/\epsilon)))$
In Table 1 we see that the randomised formulas (both CS and QF) and the QDRIFT channel all improve on the gate complexities dependence on $M$ when compared to the deterministic TS product formulas. The first order randomised TS formula (CS) scales the same as the second order deterministic formula. This is an improvement over the first order deterministic TS formula as we see an improvement in the scaling of the gate complexity with respect to the number of terms in the generator $M$, as it now scales as $M^{5/2}$ instead of $M^{3}$. We also observe that we achieve the same scaling when implementing the first order randomised formula with quantum forking. This means that there is no additional cost to performing the sampling directly on the quantum computer. The second order randomised TS (CS) gives us a quadratic dependence on $M$ which is much better than both the deterministic TS product formulas and even the first order randomised TS formula. However, as discussed in section 6.2.2 we cannot use QF to implement $S_{2}^{(ran)}$ due to the fact that we will require $2(M!)$ controlled-SWAP channels which implies that the gate complexity will depend on $M!$ which is very inefficient. For the QDRIFT channel (CS), we observe that there is no dependence on $M$ in the gate complexity. This makes this method ideal for systems with a large number of terms $M$. For example, two and three dimensional Jaynes-Cummings models with neighbour-neighbour interactions [29] and two dimensional Heisenberg models with boundary driving and local dissipation [34]. One may argue however that the $M$ dependence is hidden in the factor $\Gamma$, however for most systems we will compute $\Gamma$ classically before we construct the gate set. Therefore, this would not add a factor of $M$ to the gate complexity. For the implementation of the QDRIFT channel with QF we see that we require 2(M1)2(M-1) controlled-SWAP channels which means that the gate complexity will now scale linearly with $M$ which is still better than all other TS product formulas (deterministic and randomised). However, one should note that if we were to take into account the complexity of the best possible classical algorithms for performing the sampling on a classical computer we would still introduce an $M$ dependence, albeit additively to the complexity of the full simulation algorithm. While this scaling with $M$ is better than other methods, the quadratic scaling in $t$ is what makes this method only applicable for short simulation times, for longer simulation times, a randomised product formula will perform better.
Our approaches offer distinct advantages and trade-offs, making them suitable for different simulation scenarios, as summarized in Table 2.
Methods Pros Cons Best Use Case
Randomised TS (CS) Improves scaling with generator terms ($M$) compared to deterministic TS. Same complexity as 2nd-order deterministic TS. Implementation requires a classical co-processor for sampling. General-purpose improvement over deterministic formulas, especially for longer simulation times where scaling with t is critical.
Randomised TS (QF) Fully quantum implementation without a classical sampler. No additional complexity overhead compared to the CS version. Requires additional ancillary qubits and controlled-SWAP channels. 2nd-order version is inefficient due to M! scaling. When a ”quantum-native” implementation is preferred and the overhead of ancilla management is acceptable.
QDRIFT (CS) Gate complexity is independent of $M$. Gate complexity scales quadratically with simulation time ($t^{2}$). Systems with a very large number of generator terms ($M$) but requiring simulation for short times.
QDRIFT (QF) Quantum-native implementation. Still offers better $M$-scaling (linear) than any TS formula. Reintroduces a linear dependence on $M$ into the gate complexity. Scenarios requiring a fully quantum circuit for systems with large $M$, where a linear dependence is still a significant improvement.
Table 2: A summary of trade-offs and simulation scenarios of the methods developed in the paper
Concurrent with our work, other methods have recently been proposed that also leverage randomisation for the simulation of open quantum systems [33, 35]. For example, a qDRIFT-type approach was developed where the Lindbladian is decomposed into an ensemble of simpler generators, with analysis provided for both average and typical realizations [35]. Another approach combines a second-order product formula with randomized compiling of the dissipative dynamics, achieving improved scaling for a restricted but physically relevant class of Lindbladians [33]. Our work complements these developments by providing a direct and general extension of both the randomised Trotter-Suzuki formulas and the QDRIFT protocol to arbitrary Markovian dynamics governed by the GKSL equation. A key distinction of our analysis is the derivation of rigorous error bounds that do not rely on the mixing lemma, which is specific to Hamiltonian simulation. Furthermore, we detail two concrete implementation schemes for our protocols using both classical sampling and quantum forking.
@@ 2.1 Background
for k=2,,Mk=2,...,M. Sometimes it will be convenient to absorb the decay rates $\gamma_{k}$ into each $\mathcal{L}_{k}$, by defining $\hat{\mathcal{L}}_{k}=\gamma_{k}\mathcal{L}_{k}$ the generator can be written as
@@ 2.2 Deterministic Digital Simulation of Markovian Open Quantum Systems
The main goal of digital quantum simulation of Markovian OQS is to find novel ways of constructing an approximation $\tilde{T}_{t}$ of the total evolution $T_{t}=\exp(t\mathcal{L})$ such that for a given GKSL generator $\mathcal{L}$, a precision $\epsilon>0$, a simulation time $t\geq 0$ and a distance measure dist(,)($\cdot$$\cdot$),
@@
where $\tilde{T}_{t}$ could be implemented on a quantum computer efficiently. The most common way to obtain this approximation is by using Trotter-Suzuki (TS) product formulas [5, 6]. Using TS product formulas, the total evolution $T_{t}=\exp(t\sum_{k=1}^{M}\hat{\mathcal{L}}_{k})$ is approximated as some product of simpler channels up to a precision $\epsilon\geq 0$ when using the diamond norm as the distance measure for the quantum channels. For the purpose of this work, we define a simple channel as a fundamental building block of our simulation circuits. These include the constituent channels of the form $\exp(\tau\hat{\mathcal{L}}_{k})$ derived from the GKSL generator, as well as auxiliary operations like controlled-SWAP channels required for certain randomised implementations. The specific circuit implementation of these simple channels is generator-dependent and considered out of scope for our complexity analysis. However, for many physical systems where the generator terms $\hat{\mathcal{L}}_{k}$ are composed of local operators (e.g., Pauli strings), efficient quantum circuits for these channels can be constructed using standard unitary dilation techniques, such as the Stinespring representation [38, 39].
The same approach was introduced in [22], however, the $1\rightarrow 1$ Schatten norm was used instead of the diamond norm. We start by dividing the time $t\geq 0$ into $N\in\mathbb{N}$ steps so that we have a small time step $\tau=t/N$. Next we construct approximations of the total channel $T_{t}$ using TS product formulas that approximate the channel for a small time step $T_{\tau}$ and then simulate the TS product formula $N$ times. For example, we can approximate the evolution $T_{\tau}$ up to first order by using the following deterministic TS product formula,
@@
Let $\Lambda:=\max_{k}\left\|\hat{\mathcal{L}}_{k}\right\|_{\diamond}$ and then for $N$ chosen such that $Mt\Lambda/N\leq 1$. Then,
$\left\|T_{t}-S_{1}^{(det)}\left(\frac{t}{N}\right)^{N}\right\|_{\diamond}\leq\frac{et^{2}\Lambda^{2}M^{2}}{N},$ (21)
where we choose
$\epsilon\geq\frac{et^{2}\Lambda^{2}M^{2}}{N},$ $N\geq\frac{et^{2}\Lambda^{2}M^{2}}{\epsilon}.$ (22)
The proof of Theorem 1. can be found in Appendix A
@@
Let $\Lambda:=\max_{k}\left\|\hat{\mathcal{L}}_{k}\right\|_{\diamond}$ and we choose $N$ such that $Mt\Lambda/N\leq 1$. Then,
$\left\|T_{t}-S_{2}^{(det)}\left(\frac{t}{N}\right)^{N}\right\|_{\diamond}\leq\frac{eM^{3}t^{3}\Lambda^{3}}{3N^{2}},$ (24)
where we choose
$\epsilon\geq\frac{eM^{3}t^{3}\Lambda^{3}}{3N^{2}},$ $N\geq\frac{e^{1/2}M^{3/2}t^{3/2}\Lambda^{3/2}}{\sqrt{3\epsilon}}.$ (25)
The proof of Theorem 2. can be found in Appendix A. Now that we have shown that we can approximate $T_{t}$ by TS product formulas, all that is left is to construct a quantum circuit that implements this product formula on a quantum computer. Since our product formula is a quantum channel, one needs to use a unitary dilation, for example the Stinespring representation of the channel [38, 39], to construct a quantum circuit. However, in this work, when we draw quantum circuits we will only show the action of the channel on the state as this keeps the diagrams concise. To make clear how the diagrams should be interpreted, we note that wires in our circuits correspond to density matrices of a subsystem and gates correspond to quantum channels, thereafter the usual rules of quantum circuits may be inferred. As an illustrative example consider the deterministic product formula $S_{2}^{(det)}(t/N)^{N}$ that approximates the total evolution $T_{t}$ up to a precision $\epsilon$. This means that the output of the quantum circuit that implements $S_{2}^{(det)}$ is a density matrix $\tilde{\rho}(t)$ that is a distance $\epsilon/2$ from the density matrix $\rho(t)$. One can easily see this by using the definition of the trace distance between states i.e. $d_{tr}(\rho(t),\tilde{\rho}(t))$ and inequality (15),
$d_{tr}(\rho(t),\tilde{\rho}(t))$ $=\frac{1}{2}\norm{\rho(t)-\tilde{\rho}(t)}_{1}$
$=\frac{1}{2}\norm{T_{t}(\rho(0))-S_{2}^{(det)}(t/N)^{N}\rho(0)}_{1}$
$\leq\frac{1}{2}\left\|T_{t}-S_{2}^{(det)}(t/N)^{N}\right\|_{\diamond}$
@@
At this point we need to analyse the gate complexity of the quantum circuits we have constructed to implement the deterministic TS product formulas. To do this, we start by defining the gate complexity of our circuits as the number of simple channels that are implemented in each quantum circuit. For the case of deterministic TS product formulas, these simple channels are just the exponentials of the form $\exp(\tau\hat{\mathcal{L}}_{k})$. For example, consider the product formula $S_{1}^{(det)}(\tau)$ in (24). This formula is the product of $M$ exponentials. If we consider the quantum circuit that implements $S_{1}^{(det)}(\tau)^{N}$ to approximate $\rho(t)$ to a precision $\epsilon$, then we have $N=\left\lceil t^{2}\Lambda^{2}M^{2}/\epsilon\right\rceil$ applications of $S_{1}^{(det)}(\tau)$ which implies that we have to implement $\left\lceil t^{2}\Lambda^{2}M^{2}/\epsilon\right\rceil M$ exponentials, where $\lceil\cdot\rceil$ denotes the ceiling function. If we denote the gate complexity for the circuit that implements $S_{1}^{(det)}$ as $g_{1}^{(det)}$ then the complexity is given by
@@ 3 First Order Randomised Trotter-Suzuki Formula
Given the generator $\mathcal{L}$, as in equation (10), of a quantum channel $T_{t}$ and some time $t\geq 0$. Define the first order randomised product formula as in equation (30). Let $\Lambda:=\max_{k}\left\|\hat{\mathcal{L}}_{k}\right\|_{\diamond}$ and choose $N$ such that $Mt\Lambda/N\leq 1$. Then,
$\left\|T_{t}-S_{1}^{(ran)}\left(\frac{t}{N}\right)^{N}\right\|_{\diamond}\leq\frac{eM^{3}t^{3}\Lambda^{3}}{3N^{2}},$ (31)
where
$\epsilon\geq\frac{eM^{3}t^{3}\Lambda^{3}}{3N^{2}},$ (32)
$N\geq\frac{e^{1/2}(t\Lambda M)^{3/2}}{\sqrt{3\epsilon}}.$ (33)
@@
For the second term in equation (3), we use the fact that the diamond norm is sub-multiplicative and $\left\|\hat{\mathcal{L}}_{i}\right\|_{\diamond}\leq\Lambda$ for all i=1,,Mi=1,...,M to show that
$\left\|\hat{\mathcal{L}}_{1}^{j_{1}}\hat{\mathcal{L}}_{2}^{j_{2}}...\hat{\mathcal{L}}_{M}^{j_{M}}\right\|_{\diamond}\leq\left\|\hat{\mathcal{L}}_{1}\right\|_{\diamond}^{j_{1}}...\left\|\hat{\mathcal{L}}_{M}\right\|_{\diamond}^{j_{M}}\leq\Lambda^{j_{1}+...+j_{M}}.$ (52)
@@
Now using Lemma B.1. from Appendix B, we can compute the restricted sums in equations (53) and (54). We then get the final bounds on these terms as